a^3+b^3=0 then log (a+b) =1/2 (log a + log b + log 3)
Answer by ikleyn(52776)

.
a^3+b^3=0 then log (a+b) =1/2 (log a + log b + log 3)
~~~~~~~~~~~~~~~~~~~~~


I don't know exactly what is written here:  it is the piece of a statement,  but not the entire,  whole and/or completed mathematical statement.

OK, I will read it in this way:

    If  a^3 + b^3 = 0  then  log(a+b) = 1/2*(log a + log b + log 3)

because it is the most natural way to decipher the original text.


But if we are in the real numbers domain,  then   a^3 + b^3 = 0 implies

a^3 = -b^3,   which implies   a = - b.

Then  log (a+b) = log(0),   which is not defined.


The resume from my analysis is this:   the proposed text  (presented as is and without additional comments)  is an example

of mathematically illiterate text,

and I regret about my time which I spent for nothing, analyzing this post.